Write all momentum calculations in units of MeV/c and keep 5 significant figures in all answers. The mass of the proton is m = 938.3 MeV/c2 (This is equivalent to m = 1.673 x 10-27 kg, but MeV/c2 is a more convenient mass unit to use in atomic and nuclear physics calculations.).

Calculate the Classical Momentum of a proton traveling at 0.03 c


Calculate the Percent Error between the Classical Momentum and the Relativistic Momentum at 0.03 c. Treat the Relativistic Momentum as the accepted value.Calculate the Relativistic Momentum of a proton traveling at 0.03 c.

Calculate the Classical Momentum of a proton traveling at 0.41 c.

Calculate the Relativistic Momentum of a proton traveling at 0.41 c.

Calculate the Percent Error between the Classical Momentum and the Relativistic Momentum at 0.41 c. Treat the Relativistic Momentum as the accepted value.

Calculate the Classical Momentum of a proton traveling at 0.94 c.


Calculate the Relativiistic Momentum of a proton traveling at 0.94 c.

Calculate the Percent Error between the Classical Momentum and the Relativistic Momentum at 0.94 c. Treat the Relativistic Momentum as the accepted value
.
Answer #1

Classical momentum is defined as

Pc = m*v

Relativistic momentum is defined as

Pr = m*v /SQRT ( 1- v2/c2)

Percentage Error can be calculated as

Error = (Pr - Pc) / Pr * 100%

We have used all above expression in excel file and calculated values in below table:

Proton Mass Proton Speed|Classical Momentum Relativistic Momentum Percent (%) (in MeV/c2) (inc) n Mev/c (in MeV/c) Error 938.3 938.3 938.3 0.03 0.41 0.94 28.14900 384.70300 882.00200 28.16168 421.78391 2585.19365 0.04501 8.79145 65.88256